OpenStudy (anonymous):

calculate the derivative by expanding or simplifying the function: s(t) = (1-2t)/(t^1/2)

zepdrix (zepdrix):

\[\Large\rm s(t)=\frac{1-2t}{t^{1/2}}=\frac{1}{t^{1/2}}-\frac{2t}{t^{1/2}}\]Then we can apply some exponent rules from there.

zepdrix (zepdrix):

Remember your exponent rules? :o \[\Large\rm \frac{x^a}{x^b}=x^{a-b}\]

OpenStudy (anonymous):

i need it done by the Quotient rule (vu'-uv')/(v^2)